The increase in the volume of 10cm³ of mercury when temperature rises by 100⁰c is 0.182cm³, what is the cubic expansivity of mercury


1
Expert's answer
2021-02-19T10:31:14-0500

By definition, the increase in volume will be:




"\Delta V = \beta V_0\Delta T"

where "\beta" is the cubic expansivity, "V_0 = 10cm^3" is the initial volume, and "\Delta T = 100\degree C"is the change in temperature. Since "\Delta V = 0.182cm^3", the cubic expansivity is:



"\beta = \dfrac{\Delta V}{V_0\Delta T}\\
\beta = \dfrac{0.182}{10\cdot 100} = 1.82\times 10^{-4}\degree C^{-1}"

Answer. "1.82\times 10^{-4}\degree C^{-1}".